# Model-free Legibility: Enhancing Human-Robot Interactions through Implicit Communication and Influence Modulation

### v1

@@ Abstract
Communication is essential for successful interaction. In human-robot interaction, implicit communication enhances robots’ understanding of human needs, emotions, and intentions. This paper introduces a method to foster implicit communication in HRI without explicitly modeling human intentions or relying on pre-existing knowledge. Leveraging Transfer Entropy, we modulate influence between agents in social interactions in scenarios involving either collaboration or competition. By integrating influence into agents’ rewards within a partially observable Markov decision process, we demonstrate that boosting influence enhances collaboration or competition performance, while resisting influence diminishes performance. Our findings are validated through simulations and real-world experiments with human participants.
Keywords: Human-Robot Interaction, Implicit communication, Influence modulation
1 Introduction
@@
It is widely recognised that communication is key to successful interaction. Humans communicate with each other through both explicit (direct, deliberate communication over an established channel with clear intent to reach a defined recipient [1]) and implicit channels. Implicit communication is a subtle, indirect mode of conveying information, often relying on context, nonverbal cues, and shared understanding between communicators to convey meaning without explicit verbalization [2]. Implicit communication is particularly crucial for human-robot interaction as it enhances a robot’s ability to proactively understand and respond to human needs, emotions, and intentions, thereby facilitating more natural and effective communication and collaboration between humans and robots. Unfortunately, most current human-robot interaction (HRI) studies focusing on implicit communication explicitly model the intention of human participants [3] [4], or rely on existing intention knowledge [5]. This information is challenging to obtain in general settings.
@@
This paper proposes a method to facilitate implicit communication without the need to explicitly model human participants or rely on pre-existing knowledge. Our approach conceptualizes communication as the degree of influence agents have on one another, employing information-theoretic techniques. Specifically, we use Transfer Entropy (TE) [6] to modulate influence in social interactions. We illustrate our method using a social navigation setting [7], targeting scenarios where the objectives of individual agents may result in a need for either collaboration or competition. In particular, we demonstrate the proposed method in a corridor dilemma setting, a popular scenario in HRI [8] [9]. For our experiments, we define this as a scenario where individuals encounter each other in a narrow passage and where a given agent may wish to meet or avoid another agent. This scenario involves implicit information exchange, where participants attempt to discern each other’s intentions without verbal communication. By observing and reacting to subtle cues such as changes in trajectories, participants estimate each other’s objectives to inform their own actions. Interactions of this form can be complex, especially when both participants act simultaneously, leading to a state of ”resonance” that dissipates as mutual understanding is achieved.
@@
We consider a partially observable Markov decision process (POMDP) setting and augment the rewards of agents to include the TE that boosts or resists influence between agents. We find that boosting influence enhances the other participant’s ability to collaborate or compete with the ego-agent, while resisting influence mainly diminishes their ability to collaborate. These findings hold both in self-play [10] simulations and in real-world experiments with human participants.
2 Related Work
@@
Implicit communication has attracted significant attention in the field of robotics. Che et al. [3] present a planning framework that integrates implicit and explicit communication during mobile robot navigation, using a human behavior model and inverse reinforcement learning to generate communicative actions that enhance robot transparency and efficiency. Li et al. [11] [12] introduce a novel interaction framework that enables communication between humans and assistive devices, leveraging gaze movements to infer human intent. Liang et al. [13] investigate the role of implicature, a form of implicit communication, in HRI using the cooperative card game Hanabi, to demonstrate that human players paired with an implicature AI are significantly more likely to perceive their partner as human. These implicit communication studies typically involve explicit modeling of other participants or rely on pre-existing social knowledge [5]. In contrast, our approach is model-free, and does not require pre-exisiting social knowledge.
@@
Implicit communication is closely related to intent communication, which aims to allow observers to easily deduce the intent behind generated behaviors [14] [15]. Motions conveying intent have been shown to enhance perceived safety in virtual human-robot path-crossing tasks [16]. In reinforcement learning settings, robustness, efficiency and energy reward terms have been shown to enhance the ability of humans to interpret a robot’s intent [14]. Research focusing on projected visual legibility cues [17] has indicated that projected arrows generally offer greater interpretability than flashing lights in navigation settings. Notably, the means through which people convey and comprehend intent vary, and there exists no universally applicable method for assessing intent communication. By viewing communication as the exchange of information between agents, our TE-based framework holds the potential to adapt to various forms of communication, offering a versatile method to shape communication dynamics. Importantly, unlike many of the uni-directional intent communication approaches above, our framework captures the multi-directional nature of communication.
@@
Transfer Entropy is a measure that allows the analysis of the information transfer and potential causal relationships between two simultaneous time series. TE is most prevalent in economics literature [18, 19], but has also been used to analyse animal-animal or animal-robot interactions  [20][21], joint attention [22] and to model pedestrian evacuation [23]. Berger et al. apply TE to detect human-to-robot perturbations using low-cost sensors [24]. TE has also been used to identify arbitrary cues from raw social interaction data between humans [25]. These works successfully quantify information transfer or the relationship between sources of information, but tend to focus on specific features or aspects of interest. Our work seeks to extend the use of TE to provide a more general approach aimed at enhancing information transfer in decision-making settings leading to improved human-robot collaboration.
@@
3 Methodology
We consider an environment with multiple agents, each possessing distinct states, actions, and rewards. These rewards can be complementary (collaborative), competitive, or unrelated. The ego-agent aims to maximize its own rewards, ideally without negatively impacting other agents. In a HRI context, the robot should exhibit behaviour perceived as altruism. A fair outcome involves promoting collaboration, yielding to humans in competitive scenarios, and allowing independent agents to achieve their objectives if there is no conflict of interest. Within this environment, the ego-agent has complete access to its own state and can observe other agents, but may lack access to their decision-making processes or complete states. Consequently, we model interaction as a partially observable Markov decision process (POMDP), where states represent system situations and actions determine transitions between these states. Observations derived from states guide decision-making despite incomplete information.
@@ 3.1 Measuring influence
$T_{Y\rightarrow X}^{(k,l)}(t)$ $=I(X_{t}:\mathbf{Y}_{t}^{(l)}|\mathbf{X}_{t}^{(k)})=H(X_{t}|\mathbf{X}_{t}^{(k)})-H(X_{t}|\mathbf{X}_{t}^{(k)},\mathbf{Y}_{t}^{(l)})$ (2)
@@
$=[-\sum_{p\in P_{a_{1,t}}^{-}}p(a_{1,t})\log_{2}p(a_{1,t})]-[-\sum_{p\in P_{a_{1,t}}^{+}}p(a_{1,t})\log_{2}p(a_{1,t})].$
@@ 3.2 Q-learning
4 Experiment and Results
We investigate the proposed approach in our corridor dilemma setting (Fig. 1), a simplified 11x5 grid-world where two players start at opposite ends of a corridor with their own randomly assigned objectives: passing or meeting. In each round, players move forward and choose to move left, straight, or right. The game ends when both agents reach the same row, resulting in either a pass or a meeting, with scores based on objectives. This game can be both competitive and collaborative, requiring anticipation, strategy, and potential collaboration.
@@ 4 Experiment and Results
We apply the proposed approach using agent coordinates as states and observations. In this specific application, the reward $r$ in Equation (1) is $r=0$ when the agent takes an action before reaching the middle row of the corridor, $r=10$ when the agent successfully achieves its objective, and $r=-10$ when it fails to achieve its objective after reaching the middle row. During the experiments, the TE reward components are normalised and scaled between 0 and 10 to match the objective rewards.
@@
4.1 Simulation
We conduct a range of simulations, putting each agent type against one another in a self-play framework where pairs of agents are simultaneously trained against each other for 30000 episodes with randomised objectives. $P1$ and $P2$ refer to the two players shown in Figure 1. For each experiment, we ran 30000 episodes and gathered results from six trials with distinct random seeds and compute their averages for presentation. More simulation details can be found in Appendix 6.1.
We report the averaged success rates for both competition and collaboration (denoted as SRCP and SRCL) to assess the agents’ capacity to compete or collaborate effectively. In this context, competition refers to scenarios where agents are assigned different objectives, while collaboration refers to situations where both agents share the same objective. These success rates are quantified as the number of successful outcomes in either competition or collaboration, divided by the total number of instances of competition or collaboration. The results are presented in Table 1.
Table 1: Final average success rate of competition (SRCP) and collaboration (SRCL) in simulation.
@@ 4.1 Simulation
In competition, the Non-TE agent in the Non-TE vs. Positive-TE experiment achieved the highest final performance with a 63.38% SRCP, with a corresponding reduction in SRCP for the paired Positive-TE agent. This is as expected, since increasing influence or legibility can lead to the self-sacrificing or altruistic behaviour described in section 3.1. This increase in influence allows the Non-TE agent to learn to satisfy its objectives more frequently both in circumstances of collaboration ($P1$ and $P2$ share the same objectives) and competition ($P1$ and $P2$ share different objectives).
@@
Note that the Positive-TE vs. Positive-TE pair reaches the highest SRCL (91.72%), and also much fairer outcomes in terms of competition (49.95% vs. 50.05%) compared to the Non-TE vs. Positive-TE pair (63.38% vs. 36.62%). In addition, the Negative-TE agent in Positive-TE vs. Negative-TE group that tries to resist influence does not achieve a higher SRCP than the Non-TE agent in the Non-TE vs. Positive-TE group. These findings suggest that engaging in either receiving or resisting influence may compromise performance, but training with Positive-TE agent is beneficial for group collaboration and contributes to the fairness of the group. This is clearly evident as all pairs featuring at least one Positive-TE agent, consistently attain better collaboration performance than those without, and results in fairer competition outcomes (SRCP) for the Positive-TE vs. Positive-TE pair.
@@
These experiments clearly show that performance in interactive settings can be affected by controlling influence in basic social interactions, indicating potential applicability to broader human-robot-interaction scenarios. Additional simulation results and analysis can be found in Appendix 6.2 and 6.3. Next, we verify the efficacy of the proposed method for agents interacting with human users.
4.2 Human-Agent Experiment
We conducted a user study in the corridor dilemma environment (See Appendix 6.4 for the interface design). Here, human participants played against three agents. Typically, individuals display ambivalence, occasionally leaning slightly towards selfishness or altruism, but not to an extreme degree. Therefore, a Non-TE agent, being ambivalent, could most accurately represent a typical human. Thus, participants are tested against three agents pre-trained under this assumption: 1. Non-TE agent ($P1$) vs. Non-TE agent ($P2$), 2. Non-TE agent ($P1$) vs. Positive-TE agent ($P2$), and 3. Non-TE agent ($P1$) vs. Negative-TE agent ($P2$). Each participant was asked to play against each of the three agents (randomly ordered) for 100 consecutive rounds with randomly generated objectives. The number of rounds was selected to provide ample opportunities for each participant to familiarise themselves with an agent and allow user performance to converge to a stable level. As part of the study, participants were told that one of the opponents could potentially be a real human player. We recruited 26 participants (11 female, 15 male, aged between 20 and 37). Their collaboration and competition performances were observed across 100 rounds of gameplay against each opponent, and subsequently averaged across all participants. These findings are illustrated in Figure 2.

### v2

# Influence-Based Reward Modulation for Implicit Communication in Human-Robot Interaction

@@ Abstract
Communication is essential for successful interaction. In human-robot interaction, implicit communication holds the potential to enhance robots’ understanding of human needs, emotions, and intentions. This paper introduces a method to foster implicit communication in HRI without explicitly modelling human intentions or relying on pre-existing knowledge. Leveraging Transfer Entropy, we modulate influence between agents in social interactions in scenarios involving either collaboration or competition. By integrating influence into agents’ rewards within a partially observable Markov decision process, we demonstrate that boosting influence enhances collaboration, while resisting influence diminishes performance. Our findings are validated through simulations and real-world experiments with human participants in social navigation settings.
I Introduction
@@
It is widely recognised that communication is key to successful interaction. Humans communicate with each other through both explicit (direct, deliberate communication over an established channel with clear intent to reach a defined recipient [1]) and implicit channels. Implicit communication is a subtle, indirect mode of conveying information, often relying on context, nonverbal cues, and shared understanding between communicators to convey meaning without explicit verbalisation [2]. Implicit communication is particularly crucial for human-robot interaction as it enhances a robot’s ability to proactively understand and respond to human needs, emotions, and intentions, thereby facilitating more natural and effective communication and collaboration between humans and robots. To date, most current human-robot interaction (HRI) studies focusing on implicit communication explicitly model the intention of human participants [3] [4], or rely on existing intention knowledge [5]. This information is challenging to obtain in general settings.
@@
This paper proposes a method to facilitate implicit communication without the need to explicitly model human participants or rely on pre-existing knowledge. Our approach conceptualises communication as the degree of influence agents have on one another, employing information-theoretic techniques. Specifically, we use Transfer Entropy (TE) [6], a measure of directed information transfer, to modulate influence in social interactions. We illustrate our method using a social navigation setting [7], targeting scenarios where the objectives of individual agents may result in a need for either collaboration or competition. In particular, we demonstrate the proposed method in a corridor dilemma setting, a popular scenario in HRI [8] [9]. For our experiments, we define this as a scenario where individuals encounter each other in a narrow passage and where a given agent may wish to meet or avoid another agent (Figure 1). This scenario involves implicit information exchange, where participants attempt to discern each other’s intentions without verbal communication. By observing and reacting to subtle cues such as changes in trajectories, participants estimate each other’s objectives to inform their own actions. Interactions of this form can be complex, especially when both participants act simultaneously, leading to a state of ”resonance” that dissipates as mutual understanding is achieved.
Figure 1: Humans rely on implicit communication and information exchange to negotiate complex social settings. This work introduces a reward augmentation framework to emulate this implicit communication behaviour. We show that by optimising robot policies to increase information flow, we can influence human behaviour and improve human-robot collaboration.
@@
To better equip robots with the ability to handle this type of setting, we consider a partially observable Markov decision process (POMDP) and propose a novel approach that augments the rewards of agents to include the TE that boosts or resists influence between agents. We find that boosting influence enhances the other participant’s ability to collaborate or compete with the ego-agent, while resisting influence mainly diminishes their ability to collaborate. These findings hold both in self-play [10] simulations and in virtual human-agent experiments. Real-world human-robot experiments with physical robots also show that boosting influence improved collaboration, but reduced competition performance. Overall, our work makes the following contributions and findings:
  • We propose a novel framework for enhancing implicit communication in social HRI through influence modulation, without requiring explicit modelling or prior knowledge of humans.
  • We evaluate the framework in both simulation and virtual human-agent interaction environments, finding that promoting information transfer promotes collaboration and benefits the human in competitive scenarios.
  • We show that this collaborative effect holds in real-world human-robot experiments, but reveal interesting discrepancies in competitive interactions.
II Related Work
@@
Implicit communication has attracted significant attention in the field of robotics. Che et al. [3] present a planning framework that integrates implicit and explicit communication during mobile robot navigation, using a human behaviour model and inverse reinforcement learning to generate communicative actions that enhance robot transparency and efficiency. Li and Zhang [11] [12] introduce a novel interaction framework that enables communication between humans and assistive devices, leveraging gaze movements to infer human intent. Liang et al. [13] investigate the role of implicature, a form of implicit communication, in HRI using the cooperative card game Hanabi, to demonstrate that human players paired with an implicature AI are significantly more likely to perceive their partner as human. These implicit communication studies typically involve explicit modelling of other participants or rely on pre-existing social knowledge [5]. In contrast, our approach is model-free, and does not require pre-existing social knowledge.
@@
Implicit communication is closely related to intent communication, which aims to allow observers to easily deduce the intent behind generated behaviours [14] [15]. Motions conveying intent have been shown to enhance perceived safety in virtual human-robot path-crossing tasks [16]. In reinforcement learning settings, robustness, efficiency and energy reward terms have been shown to enhance the ability of humans to interpret a robot’s intent [14]. Research focusing on projected visual legibility cues [17] has indicated that projected arrows generally offer greater interpretability than flashing lights in navigation settings. Notably, the means through which people convey and comprehend intent vary, and there exists no universally applicable method for assessing intent communication. By viewing communication as the exchange of information between agents, our TE-based framework holds the potential to adapt to various forms of communication, offering a versatile method to shape communication dynamics. Importantly, unlike many of the uni-directional intent communication approaches above, our framework captures the multi-directional nature of communication.
Transfer Entropy (TE) is a mathematical concept introduced by Schreiber [6] that quantifies the directional flow of information between two random processes. Formally, TE is defined as the difference between the conditional entropies of one process given the past of another and the past of itself. This is expressed mathematically as
$TE(X\rightarrow Y)$ $=H(Y_{t}|Y_{t-1},Y_{t-2},\ldots)$ (1)
$-H(Y_{t}|Y_{t-1},Y_{t-2},\ldots,X_{t-1},X_{t-2},\ldots),$
@@
where $H$ denotes entropy, and $X$ and $Y$ are the two stochastic processes under consideration. By measuring how much uncertainty about the future state of one process is reduced when the past of another process is known, TE provides a rigorous framework for assessing information transfer. This ability to capture not just correlations but also the causal influence between processes makes TE particularly useful for analysing complex interactions in diverse fields such as neuroscience, economics, and human behaviour, where understanding the directionality of information flow is crucial. TE is most prevalent in economics literature [18, 19], but has also been used to analyse animal-animal or animal-robot interactions [20][21], joint attention [22] and to model pedestrian evacuation [23]. Berger et al. [24] apply TE to detect human-to-robot perturbations using low-cost sensors. TE has also been used to identify arbitrary cues from raw social interaction data between humans [25]. These works successfully quantify information transfer or the relationship between sources of information, but tend to focus on specific features or aspects of interest. Our work seeks to extend the use of TE to provide a more general approach aimed at enhancing information transfer in decision-making settings, leading to improved human-robot collaboration. Our core assumption is that greater transparency or information transfer reduces information asymmetry [26], thereby enhancing collaboration potential. In this paper we leverage TE to control information asymmetry, and tip the balance of power in favour of or against opponents.
@@
III Methodology
We consider an environment with multiple agents, each possessing distinct states, actions, and rewards. These rewards can be complementary (collaborative), competitive, or unrelated. The ego-agent aims to maximise its own rewards, ideally without negatively impacting other agents. In a HRI context, we posit that the robot should exhibit behaviour perceived as altruism. A fair outcome involves promoting collaboration, yielding to humans in competitive scenarios, and allowing independent agents to achieve their objectives if there is no conflict of interest. Within this environment, the ego-agent has complete access to its own state and can observe other agents, but may lack access to their decision-making processes or complete states. Consequently, we model interaction as a partially observable Markov decision process (POMDP), where states represent system situations and actions determine transitions between these states. Observations derived from states guide decision-making despite incomplete information.
@@ III-A Measuring influence
$T_{Y\rightarrow X}^{(k,l)}(t)$ $=I(X_{t}:\mathbf{Y}_{t}^{(l)}|\mathbf{X}_{t}^{(k)})$ (3)
$=H(X_{t}|\mathbf{X}_{t}^{(k)})-H(X_{t}|\mathbf{X}_{t}^{(k)},\mathbf{Y}_{t}^{(l)}).$
@@
$=\bigg{[}-\sum_{p\in P_{a_{1,t}}^{-}}p(a_{1,t}|\mathbf{s}_{1,t}^{(n)})\log_{2}p(a_{1,t}|\mathbf{s}_{1,t}^{(n)})\bigg{]}$
$-\bigg{[}-\sum_{p\in P_{a_{1,t}}^{+}}p(a_{1,t}|\mathbf{s}_{1,t}^{(n)},\mathbf{o}_{2,t}^{(n)})\log_{2}p(a_{1,t}|\mathbf{s}_{1,t}^{(n)},\mathbf{o}_{2,t}^{(n)})\bigg{]}.$
@@ III-B Q-learning
III-C Connections to Legibility
The proposed TE augmentation is closely related to legibility. Dragan et al. [15] formalise legible motion as motion that enables an observer to confidently infer the correct goal configuration after observing only a snippet of the trajectory. To do so, they propose a metric to evaluate legibility in a trajectory planning scenario, with the legibility L($\xi$) of a trajectory $\xi$ defined as:
[…]
where $G^{*}$ is the actual goal across the trajectory. Trajectories are more legible if the probability $P(G^{*}|\xi)$ is higher. Consider an interactive scenario with two agents, agent 1 and agent 2. Denoting the trajectory or action of agent 1 by $\tau_{1}$ and agent 2 by $\tau_{2}$, and their goals $g_{1}$ and $g_{2}$, the legibility of agent 2 from the perspective of agent 1 can be expressed as:
$L(g_{2}|\tau_{2},\tau_{1})=P(g_{2}|\tau_{2},\tau_{1}).$
Using Bayes theorem, this probability can be expressed as:
$P(g_{2}|\tau_{2},\tau_{1})=\frac{P(\tau_{2}|g_{2},\tau_{1})P(g_{2}|\tau_{1})}{P(\tau_{2}|\tau_{1})}.$
Now consider the TE from agent 1’s trajectory to agent 2’s trajectory, and for simplification, let us assume a history length of 1:
$TE(\tau_{1}\rightarrow\tau_{2})$ $=H(\tau_{2}(t)|\tau_{2}(t-1))$ (15)
$\quad-H(\tau_{2}(t)|\tau_{2}(t-1),\tau_{1}(t-1))$
$=P(\tau_{2}(t),\tau_{2}(t-1),\tau_{1}(t-1))$
$\quad\log\Bigg{[}\frac{P(\tau_{2}(t)|\tau_{2}(t-1),\tau_{1}(t-1))}{P(\tau_{2}(t),\tau_{2}(t-1))}\Bigg{]}.$
This measures the amount of information that the past states of agent 1 ($\tau_{1}(t-1)$) provide about the future states of agent 2 ($\tau_{2}(t)$), over and above the information provided by the past states of agent 2 ($\tau_{2}(t-1)$) alone. The term
$\log{\frac{P(\tau_{2}(t)|\tau_{2}(t-1),\tau_{1}(t-1))}{P(\tau_{2}(t),\tau_{2}(t-1)}}$
captures how much additional information about $\tau_{2}(t)$ is provided by $\tau_{1}(t-1)$ over and above $\tau_{2}(t-1)$. TE extends mutual information to consider the temporal dynamics and causality between the time series [6]. Thus, TE also includes mutual information about how $\tau_{1}$ affects $\tau_{2}$. This affects both $P(\tau_{2}|g_{2},\tau_{1})$ and $P(\tau_{2}|\tau_{1})$, which consequently affects the legibility $L(g_{2}|\tau_{2},\tau_{1})$. This relationship is particularly powerful, as TE gives us an implicit mechanism to modulate legibility, which is known to be challenging to measure directly.
IV Experiment and Results
We investigate the proposed approach in our corridor dilemma setting (Fig. 2), a simplified 11x5 grid-world where two players start at opposite ends of a corridor with their own randomly assigned objectives: passing or meeting. In each round, players move forward and choose to move left, straight, or right. The game ends when both agents reach the same row, resulting in either a pass or a meeting, with scores based on predefined objectives.
This game is a mixture of zero-sum (competitive) and collaborative games, requiring anticipation, strategy, and potential collaboration. We select this setting to mirror the broad range of scenarios a robot operating in the real world may encounter. These may be zero-sum (e.g., one agent wants to meet while the other does not) at times, or collaborative (e.g., both agents want to meet or pass) at others. The robot is never aware of the human’s objective, and the human does not know the robot’s objective.
@@ IV Experiment and Results
We apply the proposed approach using agent coordinates as states and observations. In this specific application, the reward $r$ in Equation (2) is $r=0$ when the agent takes an action before reaching the middle row of the corridor, $r=10$ when the agent successfully achieves its objective, and $r=-10$ when it fails to achieve its objective after reaching the middle row. The TE reward components are normalised by min-max scaling before being rescaled to between 0 and 10 to match the objective rewards. We note that TE is generally small so the game score still dominates the reward.
@@
IV-A Simulation
We conduct a range of simulations, putting each agent type against one another in a self-play framework where pairs of agents are simultaneously trained against each other for 30000 episodes with randomised objectives. $P1$ and $P2$ refer to the two players shown in Figure 2. For each experiment, we ran 30000 episodes and gathered results from six trials with distinct random seeds and compute their averages for presentation. More simulation details can be found in Appendix VIII-A.
TABLE I: Final average success rate of competition (SRCP) and collaboration (SRCL) in simulation.
@@ IV-A Simulation
We report the averaged success rates for both competition and collaboration (denoted as SRCP and SRCL respectively) to assess the agents’ capacity to compete or collaborate effectively. In this context, competition refers to scenarios where agents are assigned different objectives, while collaboration refers to situations where both agents share the same objective. These success rates are quantified as the number of successful outcomes in either competition or collaboration, divided by the total number of instances of competition or collaboration. Specifically, the success rate is calculated as follows:
Success Rate=Number of rounds that the player wonTotal number of rounds $\text{Success Rate}=\frac{\text{Number of rounds that the player won}}{\text{Total number of rounds completed}}.$
Additionally, since each round could be either collaborative or competitive for the player, we compute the SRCL and SRCP by considering only those rounds where players were either collaborating or competing. Thus, the more rounds the players completed, the more stable the metrics became. The results are presented in Table I.
@@
In competition, the Non-TE agent in the Non-TE vs. Positive-TE experiment achieved the highest final performance with a 63.38% SRCP, with a corresponding reduction in SRCP for the paired Positive-TE agent. This is as expected, since increasing influence or legibility can lead to the self-sacrificing or altruistic behaviour described in section III-A. This increase in influence allows the Non-TE agent to learn to satisfy its objectives more frequently both in circumstances of collaboration ($P1$ and $P2$ share the same objectives) and competition ($P1$ and $P2$ share different objectives).
@@
Note that the Positive-TE vs. Positive-TE pair reaches the highest SRCL (91.72%), and also much fairer outcomes in terms of competition (49.95% vs. 50.05%) compared to the Non-TE vs. Positive-TE pair (63.38% vs. 36.62%). In addition, the Negative-TE agent in Positive-TE vs. Negative-TE group that tries to resist influence does not achieve a higher SRCP than the Non-TE agent in the Non-TE vs. Positive-TE group. These findings suggest that engaging in either receiving or resisting influence may compromise performance, but training with a Positive-TE agent is beneficial for group collaboration and contributes to the fairness of the group. This is clearly evident as all pairs featuring at least one Positive-TE agent consistently attain better collaboration performance than those without, and results in fairer competition outcomes (SRCP) for the Positive-TE vs. Positive-TE pair.
@@
These experiments clearly show that performance in interactive settings can be affected by controlling influence in basic social interactions, indicating potential applicability to broader human-robot-interaction scenarios. Additional simulation results and analysis can be found in Appendix VIII-E and VIII-F. Next, we verify the efficacy of the proposed method for virtual agents interacting with human users.
IV-B Human-Agent Experiment